A random sample of size n = 90 is obtained from a population whose size is
N = 10,000 and whose population proportion with a specified characteristic is P =0.8 Describe the sampling distribution of
Answer #1

Solution :

Given : n = 90 , N = 10000 , P = 0.8

Here,

n / N = 90 / 10000 = 0.009

Which is less than 0.05 so, finite population correction factor not necessary.

Let \hat p be the sample proportion.

\therefore The sampling distribution of \hat p is approximately normal with

mean = \mu_{\hat p} =  p = 0.8

SD = \sigma_{\hat p}   =   p(1 - p)/n  

= 0.8(1 – 0.8)/90

=  0.04216370213

SD = \sigma_{\hat p} ​​​​​​​ =  0.04216370213